### v3

# Generalizable Radio-Frequency Radiance Fields for Spatial Spectrum Synthesis

Abstract
We present GRaF, Generalizable Radio-Frequency (RF) Radiance Fields, a framework that models RF signal propagation to synthesize spatial spectra at arbitrary transmitter or receiver locations, where each spectrum measures signal power across all surrounding directions at the receiver. Unlike state-of-the-art methods that adapt vanilla Neural Radiance Fields (NeRF) to the RF domain with scene-specific training, GRaF generalizes across scenes to synthesize spectra. To enable this, we prove an interpolation theory in the RF domain: the spatial spectrum from a transmitter can be approximated using spectra from geographically proximate transmitters. Building on this theory, GRaF comprises two components: (i) a geometry-aware Transformer encoder that captures spatial correlations from neighboring transmitters to learn a scene-independent latent RF radiance field, and (ii) a neural ray tracing algorithm that estimates spectrum reception at the receiver. Experimental results demonstrate that GRaF outperforms existing methods on single-scene benchmarks and achieves state-of-the-art performance on unseen scene layouts. Our source code is publicly available at https://github.com/kangyang73/GRaF.
Figure 1: In the training scene, Radio-Frequency (RF) signals from each transmitter are measured across all surrounding directions by the receiver to form a spatial spectrum. Trained on this scene, GRaF synthesizes spectra for arbitrary transmitter locations in unseen scenes.
1 Introduction
The evolution of wireless communication systems, such as WiFi and Sixth Generation (6G) cellular networks, has enabled transformative applications in domains such as smart cities and smart healthcare [13, 38, 19, 40]. These networks function as both communication backbones and sensing platforms. Reliable communication requires meticulous network planning to mitigate issues such as dead spots and outages [1], while sensing applications, e.g., WiFi-based localization [3, 34], depend on large volumes of high-quality data to train deep neural networks (DNNs). Site surveys [15, 48, 49] are conducted to collect data for communication and sensing. However, conducting these surveys demands dense measurements at numerous locations, which is both time-consuming and labor-intensive [10, 18].
An alternative is to synthesize Radio-Frequency (RF) datasets via propagation modeling, generating spatial spectra that capture the received power from all directions around the receiver (§2.1). Given transmitter and receiver positions, RF propagation modeling estimates the received spectrum while accounting for reflection, diffraction, and scattering [17, 32]. In free space, Maxwell’s equations [28] enable accurate signal computation [45]; however, in complex real-world environments, directly solving them becomes intractable [57]. To approximate these scenarios, ray tracing simulations [54, 12, 27, 14] model propagation paths using Computer-Aided Design (CAD) scene models [23, 31]. However, obtaining accurate CAD models and the high computational cost of ray tracing make such simulations both impractical and unreliable in practice.
Recently, NeRF2 [58] and NeWRF [24] have adapted Neural Radiance Fields (NeRF) [30, 36, 21, 26], originally designed for novel view synthesis in the optical domain, to model RF signal propagation and synthesize spectrum. These methods achieve state-of-the-art performance in RF data synthesis by learning neural radiance fields tailored to RF signals. However, like vanilla NeRF, they tend to overfit the scenes they are trained on, requiring time-intensive retraining for each new scene [30, 41, 22].
This paper introduces GRaF, which enhances RF spatial spectrum synthesis within a scene and generalizes to unseen scenes. As illustrated in Figure 1, spectra collected from multiple transmitter locations in a training scene are used to train GRaF. After training, GRaF can synthesize high-quality spatial spectra for arbitrary locations in new scenes. This capability is supported by our Interpolation Theorem 1 (§4.1), which establishes that the spectrum at any transmitter location can be approximated by interpolating the spectra of geographically proximate transmitters.
Building on this theory, GRaF learns a latent RF radiance field from the spectra of neighboring transmitters, enabling spatial spectrum synthesis beyond the training scene. Constructing such a field requires a fundamentally different approach from optical radiance fields, due to the distinct propagation characteristics of RF signals. Unlike visible light, RF signals have centimeter-scale wavelengths and interact with obstacles through more complex mechanisms, including absorption, reflection, diffraction, and scattering [17, 32]. To account for these interactions, we represent voxel attributes as latent vectors in a learned radiance field, constructed using a geometry-aware Transformer encoder with cross-attention. This encoder processes the spatial spectra of neighboring transmitters along with their geometric relationships to the target transmitter location.
Furthermore, we propose a neural ray tracing algorithm that operates on the latent RF radiance field to estimate the spatial spectrum. This algorithm traces rays outward from the receiver, sampling voxels along each path to account for interactions within the scene. At each voxel, it predicts complex-valued contributions that capture both amplitude and phase variations due to RF propagation. These contributions are then aggregated along each ray, incorporating free-space path loss and phase shifts induced by propagation delay, to compute the received spatial spectrum.
The key contribution of this work is a new paradigm for generalizable RF spatial spectrum synthesis, grounded in an interpolation theory that estimates the spectrum at a target location from the spatial spectra of nearby transmitters. GRaF replaces scene-specific NeRF-style approaches with a geometry-aware Transformer that encodes spatial correlations from neighboring transmitters into a latent RF radiance field, followed by a neural ray tracing algorithm that aggregates these features to compute the spatial spectrum. We demonstrate that GRaF achieves state-of-the-art performance in both single-scene and cross-scene settings, and further enhances downstream localization performance.
2 Preliminary
(a) 3D spherical view
(b) 2D projection view
Figure 2: Illustration of the spatial spectrum in 3D and 2D views.
2.1 Spatial Spectrum
The spatial spectrum characterizes how received RF signal power is distributed across 3D directions (azimuth and elevation) around the receiver when the transmitter emits RF signals. We first present the wireless channel model and then describe how the spatial spectrum is computed.
Wireless Channel Model. A wireless communication system consists of a transmitter and a receiver. The transmitted signal can be represented as a complex number $x=Ae^{j\psi}$, where $A$ is the amplitude and $\psi$ is the phase. As the signal propagates through the wireless channel, its amplitude is attenuated by $A_{\text{att}}$, and its phase is shifted by $\Delta\psi$. The signal received at a single-antenna receiver is given by:
$y=x\cdot A_{\text{att}}e^{j\Delta\psi}=A\cdot A_{\text{att}}e^{j(\psi+\Delta\psi)}.$ (1)
For a receiver configured as a $M\times N$ antenna array, the received signal is represented as a matrix $\mathbf{Y}=\big[y_{m,n}\big]_{m,n=0}^{M-1,N-1}$, where $y_{m,n}$ denotes the signal received at the $(m,n)$-th array element. Each array element $y_{m,n}$ is computed as follows:
$y_{m,n}=A\cdot A_{\text{att}}e^{j(\psi+\Delta\psi+\Delta\sigma_{m,n})},$ (2)
where $A_{\text{att}}e^{j\Delta\psi}$ is the attenuation and phase shift introduced by the channel. The term $\Delta\sigma_{m,n}$ is the geometric phase shift at the $(m,n)$-th element, caused by its position relative to the array center at $(0,0,0)$. It is determined by the element position and the angles of arrival:
$\Delta\sigma_{m,n}=\frac{2\pi}{\lambda}\left(md\sin\beta\cos\alpha+nd\sin\beta\sin\alpha\right),$ (3)
where $\lambda$ is the signal wavelength, $d=\frac{\lambda}{2}$ is the element spacing, and $\alpha$ and $\beta$ are the azimuth and elevation angles of the incoming signal, respectively.
Spatial Spectrum Computation. It is computed from the received signal matrix $\mathbf{Y}$ to estimate the power distribution over azimuth $\alpha$ and elevation $\beta$ angles. For a given direction $(\alpha,\beta)$, the received signal power is given by:
$\mathbf{SS}(\alpha,\beta)=\left|\sum_{m=0}^{M-1}\sum_{n=0}^{N-1}y_{m,n}e^{-j\Delta\sigma_{m,n}}\right|^{2},$ (4)
where $\Delta\sigma_{m,n}$ is the theoretical geometric phase shift defined in Equation (3). With one-degree resolution and only the upper hemisphere considered, the spatial spectrum $\mathbf{SS}$ is represented as a matrix of size $(N_{a},N_{e})$: $\mathbf{SS}=\big[SS_{p,q}\big]_{p=0,q=0}^{N_{a}-1,N_{e}-1},$ where $N_{a}=360$ (azimuth) and $N_{e}=90$ (elevation). Figure 2(a) shows the 3D spatial spectrum, while Figure 2(b) depicts the corresponding 2D projection onto the X-Y plane, where the radial distance is $\cos(\beta)$.
2.2 Vanilla NeRF for Spatial Spectrum Synthesis
Vanilla NeRF has recently been adapted for RF spatial spectrum synthesis [58, 24]. It models a 3D scene using a Multi-Layer Perceptron (MLP) with 8-dimensional inputs:
$\mathcal{F}_{\Theta}:\left(\{x,y,z\},\alpha,\beta,\{x_{tx},y_{tx},z_{tx}\}\right)\to\left(\delta,\xi\right),$ (5)
where $\{x,y,z\}$ are the coordinates of a voxel, and $(\alpha,\beta)$ denote the azimuth and elevation angles defining the view direction of a ray traced from the receiver. The transmitter coordinates $\{x_{tx},y_{tx},z_{tx}\}$ are also included, as each voxel’s properties are influenced by the transmitter’s location. The outputs $(\delta,\xi)$ represent the voxel’s RF characteristics, where $\delta$ is the attenuation and $\xi$ is the signal emission, with each voxel treated as a secondary RF signal source [5].
To generate spatial spectrum, NeRF2 [58] performs ray tracing in each direction of the spectrum. Discrete sample voxels $\{V_{1},\ldots,V_{S}\}$ are taken along each ray $\mathbf{r}$, and the MLP is queried with their coordinates to predict $\delta$ and $\epsilon$. The result for each ray, $\mathbf{SS}(\mathbf{r})$, is calculated as:
$\mathbf{SS}\left(\mathbf{r}\right)=\sum_{i=1}^{S}\exp\left(\sum_{j=1}^{i-1}\delta_{j}\right)\xi_{i},$ (6)
where RF signals re-emitted by voxels along the ray are aggregated, with each voxel acting as a source whose contribution is attenuated by intervening voxels.
3 Related Work
Generalization in Optical NeRFs. Optical NeRFs have revolutionized novel view synthesis by learning scene representations from images. However, vanilla NeRF requires extensive per-scene training, limiting its generalization to new scenes [30, 41, 36]. Generalizable NeRFs overcome this limitation by adapting to unseen scenes without retraining [22, 6, 39, 46, 9, 37, 52, 7, 59, 4, 47]. For example, MVSNeRF integrates multi-view stereo with neural rendering to reconstruct radiance fields from a few input views [6], and WaveNeRF employs wavelet-based representations for generalization [46]. GSNeRF incorporates semantics to generate novel views and semantic maps for unseen scenes [9]. These methods cannot be directly applied to RF signals due to fundamental differences in wavelength and propagation behaviors. This work extends generalizable NeRFs to RF domain for spatial spectrum synthesis.
RF Spatial Spectrum Synthesis. Several recent works have explored NeRF-style methods in the RF domain [58, 33]. For example, NeRF2 [58] trains an MLP to compute voxel attributes and uses Equation (6) for ray tracing to compute the spatial spectrum. However, it requires training a separate model for each scene, limiting scalability and hindering generalization to new scenes. NeWRF [24] leverages direction-of-arrival (DoA) priors to reduce ray sampling and improve efficiency, but it requires specialized antenna arrays for DoA measurements, which are often impractical. WiNeRT [33] employs differentiable ray tracing with CAD-based scene geometry, but accurate CAD models are difficult to obtain in practice. RFScape [8] models RF propagation using neural Signed Distance Function (SDF)–based scene geometry, but it requires known object shapes and layouts, limiting practicality. In contrast, GRaF requires no scene-specific training, no DoA measurements, and no prior scene models, making it a practical solution for RF spatial spectrum synthesis. GeRaF [25] reconstructs 3D object geometry from near-field RF scans, whereas our work models scene-level RF spectrum synthesis and generalizes across different scenes.
3DGS [16, 53, 51]-based spectrum synthesis methods [55, 43, 50, 56, 20] improve the training and inference times of NeRF2 [58] within a single scene. In contrast, our work focuses on achieving spatial generalization across scenes.
4 Methodology
Figure 3: Architecture of GRaF. Each voxel is represented by a feature $\mathbf{v}^{s}_{i}$, where $i$ indexes the $M$ rays, and $s$ denotes the voxel’s position along the $i$-th ray. A neural-driven ray tracing algorithm computes the received signal power for each ray (TX: transmitter, RX: receiver).
Given a set of training scenes $\{1,\dots,K\}$, each containing spatial spectra and transmitter positions $\mathcal{S}^{(k)}=\left\{\big(\mathbf{SS}_{i}^{(k)},\mathbf{P}_{i}^{(k)}\big)\right\}_{i=1}^{N_{k}},$ where $\mathbf{SS}_{i}^{(k)}\in\mathbb{R}^{360\times 90}$ denotes the spatial spectrum and $\mathbf{P}_{i}^{(k)}\in\mathbb{R}^{3}$ is the transmitter position, the objective is to learn a model $\mathcal{F}_{\Theta}$ that synthesizes the spectrum $\mathbf{SS}^{(j)}_{\text{target}}$ for a transmitter at position $\mathbf{P}^{(j)}_{\text{target}}$ in any scene $j$, using its $L$ nearest neighboring transmitters $\mathcal{N}_{L}^{(j)}\subset\mathcal{S}^{(j)}$. Justification is given in §F.1 of the Supplementary Materials. This objective can be formulated as follows:
$\Theta^{*}=\arg\max_{\Theta}\prod_{k=1}^{K}\prod_{i=1}^{N_{k}}p\!\left(\mathbf{SS}_{i}^{(k)}\,\middle|\,\mathcal{N}_{L}^{(k)}\!\left(\mathbf{P}_{i}^{(k)}\right),\mathbf{P}_{i}^{(k)},\Theta\right),$ (7)
where $\mathcal{N}_{L}^{(k)}\!\left(\mathbf{P}_{i}^{(k)}\right)$ denotes the $L$ nearest transmitters to $\mathbf{P}_{i}^{(k)}$ in $\mathcal{S}^{(k)}$. At inference time, for any scene $j$ (seen or unseen during training), given a target transmitter position $\mathbf{P}_{\text{target}}^{(j)}$ and its neighbors $\mathcal{N}_{L}^{(j)}\subset\mathcal{S}^{(j)}$, the learned model synthesizes the corresponding spatial spectrum as:
$\mathbf{SS}_{\text{target}}^{(j)}=\mathcal{F}_{\Theta^{*}}\!\left(\mathcal{N}_{L}^{(j)}\!\left(\mathbf{P}_{\text{target}}^{(j)}\right),\mathbf{P}_{\text{target}}^{(j)}\right).$ (8)
4.1 RF Spatial Spectrum Interpolation Theory
Theorem 1 (Spectrum Interpolation).
Let $\mathbf{SS}(\mathbf{P})$ denote the spatial spectrum of a transmitter at position $\mathbf{P}\in\mathbb{R}^{3}$, and let $\mathcal{N}_{L}(\mathbf{P})=\{\mathbf{P}_{i}\}_{i=1}^{L}$ be its $L$ nearest neighbors with spectra $\mathbf{SS}_{i}$. Then the spectrum at $\mathbf{P}$ admits the approximation $\mathbf{SS}(\mathbf{P})\approx\sum_{i=1}^{L}w_{i}\,\mathbf{SS}_{i}$, for barycentric weights $\{w_{i}\}$ determined by the local transmitter geometry. The interpolation error satisfies $\epsilon\leq K\,\delta^{2}$, where $\delta=\max_{i}\|\mathbf{P}-\mathbf{P}_{i}\|$ denotes the neighborhood radius and $K$ characterizes the environment curvature. The full theorem and proof are provided in §A of the Supplementary Materials.
4.2 Model Overview
As illustrated in Figure 3, GRaF comprises two key components: (i) a latent RF radiance field learned by a geometry-aware Transformer that captures spatial correlations from neighboring transmitters, and (ii) a neural ray tracing algorithm that aggregates latent voxel features along directional rays to compute the spatial spectrum.
Central to this process is a latent variable $\mathbf{Z}$ [29] that encapsulates the scene’s RF propagation characteristics, such as path loss and multipath effects. In particular, $\mathbf{Z}$ represents the interpolation weights from Theorem 1, refined through non-linear transformations to capture propagation behaviors beyond simple linear combinations. Accordingly, the likelihood of the spatial spectrum $\mathbf{SS}$ given the $L$ nearest neighbors $\mathcal{N}_{L}$ and transmitter position $\mathbf{P}$ is given by:
$p\left(\mathbf{SS}\,\middle|\,\mathcal{N}_{L},\mathbf{P}\right)=\int p\left(\mathbf{SS}\,\middle|\,\mathbf{Z},\mathbf{P}\right)\,p\left(\mathbf{Z}\,\middle|\,\mathcal{N}_{L}\right)\,d\mathbf{Z},$ (9)
where $\mathcal{N}_{L}=\mathcal{N}_{L}(\mathbf{P})$ denotes the $L$ nearest transmitters to $\mathbf{P}$. Since $\mathbf{Z}$ encodes global RF propagation properties (e.g., materials and geometry), the signal power along each ray depends only on $\mathbf{Z}$. The neural ray tracing algorithm applies the interpolation theory directionally by deriving ray-specific weights from $\mathbf{Z}$ to estimate the spectrum for each direction. Thus, the likelihood decomposes across rays as:
$p\left(\mathbf{SS}\,\middle|\,\mathcal{N}_{L},\mathbf{P}\right)=\int\prod_{r=1}^{Q}p\left(\mathbf{SS}(r)\,\middle|\,\mathbf{Z},\mathbf{P}\right)\,p\left(\mathbf{Z}\,\middle|\,\mathcal{N}_{L}\right)\,d\mathbf{Z},$ (10)
where $Q=N_{a}\times N_{e}$ denotes the total number of rays discretized over azimuth $(\alpha)$ and elevation $(\beta)$ angles. For example, $Q=360\times 90$ at one-degree resolution, covering the upper hemisphere centered at the receiver.
4.3 Latent RF Radiance Field
The latent RF radiance field computes the latent variable $\mathbf{Z}$, which encodes the scene’s RF propagation characteristics, including path loss, shadowing, and multipath interactions (e.g., reflection and diffraction). This contextual representation enables GRaF to generalize spectrum synthesis across diverse scenes by providing a unified encoding of RF propagation behavior. The latent variable $\mathbf{Z}$ is obtained using a parameterized function $\mathcal{T}_{\Psi}$ that takes as input the $L$ nearest neighbors $\mathcal{N}_{L}$ and the transmitter position $\mathbf{P}\in\mathbb{R}^{3}$:
$\mathbf{Z}=\mathcal{T}_{\Psi}\left(\mathcal{N}_{L},\mathbf{P}\right).$ (11)
The architecture of $\mathcal{T}_{\Psi}$ is designed to effectively integrate both spectral and geometric information from $\mathcal{N}_{L}$. It begins with a ResNet-18 feature extractor that processes each neighbor’s spatial spectrum $\mathbf{SS}_{i}\in\mathcal{N}_{L}$, producing a compact feature vector that captures high-level patterns such as directional power distribution and signal strength variations. In parallel, the relative positions $(\mathbf{P}_{i}-\mathbf{P})$ (for each neighbor $\mathbf{P}_{i}\in\mathcal{N}_{L}$) are encoded using positional embeddings to preserve the spatial relationships between the target transmitter and its neighbors. These spectral features and geometric embeddings capture both the RF signal characteristics and the geometric context of the scene.
The joint spectral–geometric representations are then processed by a geometry-aware Transformer encoder equipped with cross-attention layers. The cross-attention mechanism dynamically emphasizes the most influential neighboring transmitters by weighting their spectral and geometric features, effectively learning the interpolation weights described in Theorem 1 while capturing non-linear propagation effects such as interference, diffraction, and multipath scattering. The output of the Transformer encoder is subsequently refined by two MLPs, which map the intermediate representation to the final latent variable $\mathbf{Z}\in\mathbb{R}^{d}$. The resulting $\mathbf{Z}$ serves as a compact contextual encoding that summarizes the scene’s RF propagation behavior and the spatial relationships among neighboring transmitters, enabling GRaF to synthesize spatial spectra.
4.4 Neural Ray Tracing Algorithm
The neural ray tracing algorithm synthesizes the spatial spectrum $\mathbf{SS}_{\Theta}$ by leveraging the contextual latent variable $\mathbf{Z}$. It traces directional rays from the receiver and aggregates latent features along each ray to estimate the signal power in that direction, producing the complete spatial spectrum.
For each ray $r$ with direction $(\alpha,\beta)$, we sample $S$ points $\{\mathbf{x}_{s}\}_{s=1}^{S}$ along its path, evenly spaced to cover the spatial extent of the scene. At each sampled point $\mathbf{x}_{s}$, a voxel-specific feature $\mathbf{v}_{s}\in\mathbb{R}^{d}$ is computed by combining the latent variable $\mathbf{Z}$, the transmitter position $\mathbf{P}\in\mathbb{R}^{3}$, and positional encodings of $\mathbf{x}_{s}$ and the ray direction $(\alpha,\beta)$:
$\mathbf{v}_{s}=\mathrm{MLP}\!\left(\mathbf{Z},\;\mathrm{PosEnc}\!\left(\mathbf{x}_{s},(\alpha,\beta),\mathbf{P}\right)\right),$ (12)
where $\mathrm{MLP}$ is a multi-layer perceptron and $\mathrm{PosEnc}$ denotes the positional encoding function. The resulting features are then processed by an RF-related function $\mathcal{R}_{\Phi}$ to estimate the signal power for the ray:
$\hat{\mathbf{SS}}_{\Theta}(r)=\mathcal{R}_{\Phi}\left(\{\mathbf{v}_{s}\}_{s=1}^{S}\right).$ (13)
To incorporate RF-specific characteristics, $\mathcal{R}_{\Phi}$ models the radiated complex signal and the complex-valued attenuation at each sampled point $\mathbf{x}_{s}$ along the ray. Inspired by the wireless channel model, the radiated signal $s(\mathbf{x}_{s},\alpha,\beta)$ and attenuation $a(\mathbf{x}_{s},\alpha,\beta)$ are computed as:
$s(\mathbf{x}_{s},\alpha,\beta)$ $=I_{s}(\mathbf{x}_{s},\alpha,\beta)+j\,Q_{s}(\mathbf{x}_{s},\alpha,\beta)$ (14)
$=\mathrm{MLP}_{s}(\mathbf{v}_{s}),$
$a(\mathbf{x}_{s},\alpha,\beta)$ $=A_{s}(\mathbf{x}_{s},\alpha,\beta)+j\,B_{s}(\mathbf{x}_{s},\alpha,\beta)$
$=\mathrm{MLP}_{a}(\mathbf{v}_{s}),$
where $\mathrm{MLP}_{s}$ and $\mathrm{MLP}_{a}$ map the feature $\mathbf{v}_{s}$ to the complex signal and complex attenuation, respectively. The attenuation term $a(\mathbf{x}_{s},\alpha,\beta)$ encodes both amplitude reduction and phase shift introduced at $\mathbf{x}_{s}$. The received ray signal $y_{r}$ is then obtained by aggregating the contributions from all sampled points along the ray. A detailed derivation and proof are provided in §B of the Supplementary Materials.
$y_{r}=\sum_{s=1}^{S}\left(\prod_{j=1}^{s-1}a\left(\mathbf{x}_{j},\alpha,\beta\right)\right)s\left(\mathbf{x}_{s},\alpha,\beta\right)\cdot\frac{\lambda}{4\pi d_{s}}e^{-j\frac{2\pi fd_{s}}{c}},$ (15)
this aggregation process can be modeled using a Transformer by mapping voxel-wise features $(\delta,\xi)$ into token features, where the attention scores naturally represent cumulative attenuation along the ray. Finally, the predicted spatial spectrum value for ray $r$ is obtained as the power of its received signal: $\hat{\mathbf{SS}}_{\Theta}(r)=|y_{r}|^{2}.$
This procedure is applied to all rays, which discretize the spectrum at one-degree resolution over azimuth and elevation angles, resulting in the complete spatial spectrum $\mathbf{SS}_{\Theta}$.
Loss Function. With the spectrum $\mathbf{SS}_{\Theta}$ computed, the overall parameters $\Theta$, which include $\Psi$ (from $\mathcal{T}_{\Psi}$) and $\Phi$ (from $\mathcal{R}_{\Phi}$), are optimized to maximize the log-likelihood of the observed spatial spectra, as defined in Equation (7):
$\Theta^{*}$ $=\arg\max_{\Theta}\log p\left(\mathbf{SS}_{\Theta}\mid\mathcal{N}_{L},\mathbf{P}\right)$ (16)
$=\arg\max_{\Theta}\log\int\prod_{r=1}^{Q}p\left(\mathbf{SS}(r)\mid\mathbf{Z},\mathbf{P}\right)\,p\left(\mathbf{Z}\mid\mathcal{N}_{L}\right)\,d\mathbf{Z}.$
Direct computation of this integral is intractable due to the high dimensionality of $\mathbf{Z}$. Because $\mathbf{Z}$ is deterministically produced by the latent RF radiance field, the optimization reduces to a supervised reconstruction problem. The derivation is provided in §C of the Supplementary Materials.
$\Theta^{*}=\arg\min_{\Theta}\sum_{r=1}^{Q}\left\|\mathbf{SS}(r)-\hat{\mathbf{SS}}_{\Theta}(r)\right\|^{2}.$ (17)
5 Experiments
(a) Conference room
(b) Office room
Figure 4: Top-view visualization of two room layouts.
Implementation details are provided in §D of the Supplementary Materials. We first evaluate GRaF in a single-scene setting, and then assess its cross-scene performance.
5.1 Experimental Setup
• Datasets (i) RFID Dataset: Introduced by NeRF2 [58], RF Identification (RFID) operates in 915 MHz, with a receiver positioned at a fixed location and equipped with a $4\times 4$ antenna array. The dataset consists of 6,123 transmitter locations, each associated with a spectrum.
(ii) MATLAB Dataset: We consider two representative indoor layouts: a conference room and an office room. Their geometries are illustrated in Figure 4, with dimensions of $9.8\,\text{ft}\times 9.8\,\text{ft}\times 8.2\,\text{ft}$ and $26.2\,\text{ft}\times 16.4\,\text{ft}\times 9.8\,\text{ft}$, respectively. Each layout is represented by a CAD model and further diversified into multiple scene versions by modifying object placements using Blender. For the conference room layout (Figure 4(a)), we define three scene versions: ConferenceV1 includes both Chairs 1 and 2, ConferenceV2 includes only Chair 1, and ConferenceV3 includes neither Chair 1 nor Chair 2. For the office room layout (Figure 4(b)), we similarly define three scene versions: OfficeV1 includes both Tables 1 and 2, OfficeV2 includes only Table 1, and OfficeV3 includes neither Table 1 nor Table 2.
In each scene version, a receiver with a fixed antenna array is positioned, and transmitters are randomly distributed throughout the space. A total of 4,416, 4,453, 3,107, 8,481, 7,274, and 4,894 transmitters are placed to collect spatial spectra for each scene version using MATLAB ray tracing simulation [27]. Unless stated otherwise, 80% of the data is for training and 20% for testing.
(iii) Additional layouts, i.e., a bedroom and an outdoor scene, are described in §E of the Supplementary Materials.
Table 1: Model performance for the single-scene settings.
[…]
KNN 0.089 0.357 15.16 0.543
KNN-DL 0.048 0.198 20.81 0.675
NeRF2 0.052 0.274 19.93 0.704
GRaF 0.038 0.136 21.94 0.766
• Baselines We compare GRaF against three baselines.
(i) K-Nearest Neighbors (KNN): It predicts the spatial spectrum at a target transmitter location by averaging the spectra of the $L$ nearest neighbors, ensuring $L$ matches GRaF for a fair comparison.
(ii) KNN-DL: Similar to KNN, this approach assigns a weight matrix of dimension $(N_{a},N_{e})$ to each neighbor instead of using equal averaging. Each pixel of the spectrum is given a learnable weight, and the target spectrum is obtained by a weighted summation of all neighbors’ spectra using their corresponding weight matrices. The weight matrices are optimized using an $\ell_{2}$ loss. Further details are provided in §A.2 of the Supplementary Materials.
(iii) NeRF2: It is introduced in §2.2. NeWRF [24] is essentially the same as NeRF2, but additionally uses DoA measurements to reduce the number of rays. Since DoA data is difficult to obtain in practice, we treat NeRF2 and NeWRF as equivalent methods in our evaluation. Other methods, e.g., WiNeRT [33] and RFScape [8], are not directly comparable since they require prior scene models.
[…]
5.2 Single-Scene Performance
There are seven scenes in total: one from the RFID dataset and six from the MATLAB dataset (ConferenceV1–V3, OfficeV1–V3). In this section, we train and evaluate a separate model for each scene, using 80% of the data for training and the remaining 20% for testing.
Figure 5: Visual comparison of synthesized spatial spectra.
Analysis. We first present the spatial spectra of four randomly selected transmitter locations, alongside the spectra synthesized by NeRF2 and GRaF, as illustrated in Figure 5. Visually, the spectra synthesized by GRaF closely match the ground truth, outperforming those generated by NeRF2. Consistent with this visual comparison, GRaF achieves superior metric scores, as illustrated in Table 1. Compared to the KNN and KNN-DL, GRaF improves PSNR by 44.7% and 5.5%, respectively. These results demonstrate that neighboring spatial spectra are indeed helpful for predicting the target transmitter’s spectrum. However, KNN’s simple averaging approach fails to capture the intricate spatial relationships among the neighboring spectra. KNN-DL further improves upon KNN’s performance, indicating that a careful interpretation of neighboring spectra can enhance the prediction accuracy. Nevertheless, KNN-DL does not account for the geometric relationships among spectra, limiting its effectiveness. Finally, GRaF outperforms NeRF2 by 26.9%, 50.4%, 10.2%, and 8.8% in terms of MSE, LPIPS, PSNR, and SSIM, respectively. This improvement highlights GRaF’s ability to leverage neighboring spatial spectra to learn latent voxel features that capture complex propagation behaviors and dynamically assign weights using an attention mechanism during ray tracing.
5.3 Generalization to Unseen Scenes
We evaluate the generalization capability of GRaF across unseen scenes, particularly when object configurations within the layout are changed. Specifically, we train two models separately for the conference room and office room layouts. For the conference room, the model is trained on ConferenceV1 and tested on the test splits of ConferenceV2 and ConferenceV3. For the office room, the model is trained on OfficeV1 and tested on OfficeV2 and OfficeV3. All baseline methods are trained and evaluated under the same settings for fair comparison. The inverse setting, i.e., the training and testing scenes are swapped, is provided in §E.2 of the Supplementary Materials.
Analysis. Table 2 presents the average quantitative results for the two layout models. Compared to the single-scene setting (Table 1), KNN achieves similar performance in both scenarios, highlighting that neighboring spatial information consistently provides value. In contrast, the performance of KNN-DL and NeRF2 declines in the unseen scene setting due to their reliance on models overfitted to the training scenes, which limits their generalization capabilities. However, because the differences between the training and testing scenes involve only minor modifications, such as adding or removing one or two tables or chairs, the performance degradation is moderate, with PSNR drops of $8.51\%$ for KNN-DL and $12.90\%$ for NeRF2. By comparison, GRaF achieves the highest accuracy in both settings by leveraging geometry-aware latent RF radiance fields and a neural ray tracing algorithm, enabling effective generalization across diverse scene layouts.
Table 2: Model performance on the unseen scenes settings.
[…]
KNN 0.083 0.361 14.77 0.552
KNN-DL 0.053 0.279 19.04 0.614
NeRF2 0.065 0.337 17.36 0.691
GRaF 0.039 0.215 20.96 0.705
5.4 Generalization to Unseen Layouts
We further evaluate the generalization capability of GRaF under a more challenging setting. Specifically, a model is trained on all three versions of the conference layout, i.e., the training splits of ConferenceV1–V3, and tested on all three versions of the office layout, i.e., OfficeV1–V3. The reverse experiment is also conducted, where the model is trained on OfficeV1–V3 and tested on ConferenceV1–V3. For fair comparison, the three baseline methods are evaluated under the same experimental settings.
Analysis. Table 3 presents the quantitative results for unseen layouts. KNN demonstrates consistent performance across settings with slight PSNR fluctuations, reflecting the effectiveness of neighboring spatial spectrum information. In contrast, KNN-DL and NeRF2 show significant performance declines in unseen layouts compared to single-scene settings. KNN-DL’s PSNR drops by $28.2\%$ and NeRF2’s by $35.9\%$, highlighting their limited generalization due to scene-specific training. GRaF achieves the highest accuracy across all settings, demonstrating robust generalization. This is attributed to its geometry-aware voxel features and neural ray tracing, which enable it to adapt to diverse scenes. However, the spectra synthesized by GRaF in cross-layout experiments are of lower quality than those in single-scene experiments. This performance decline is due to challenges posed by varied layouts and object materials [45].
The trained GRaF is also evaluated on additional layouts (Bedroom and Outdoor) as well as a real-world dataset, introducing diverse spatial characteristics. The Bedroom features confined spaces with rich multipath propagation; the Outdoor represents a large-scale open environment with sparse reflectors; and the RFID dataset introduces real-world noise, hardware variability, and temporal effects. Results are provided in §E.3 of the Supplementary Materials.
Table 3: Model performance on the unseen layouts settings.
[…]
KNN 0.085 0.359 15.32 0.539
KNN-DL 0.087 0.429 14.94 0.498
NeRF2 0.092 0.477 12.76 0.481
GRaF 0.042 0.268 17.81 0.629
Table 4: Ablated versions of GRaF in unseen scene settings.
[…]
Without cross-attention layer 0.239 19.37
Without neural ray tracing 0.379 16.79
Full model (GRaF) 0.215 20.96
5.5 Ablation Study
This section follows the unseen-scene setting described in §5.3, with results shown in Table 4. The impact of neighbor number $L$ is given in §E.4 of the Supplementary Materials.
Latent RF Radiance Field. In the parameterized function $\mathcal{T}_{\Psi}$ of our model, we adopt a geometry-aware cross-attention layer to learn contextual latent representations. To evaluate the effect of attention mechanisms, we replace the cross-attention layer with a simpler dot-product attention layer. A comparison between the first and third rows in Table 4 shows that GRaF achieves slightly better performance with the cross-attention mechanism. This result suggests that GRaF ’s generalization ability stems from its capacity to model interactions across neighboring spatial spectra.
Neural Ray Tracing. We replace our neural ray tracing algorithm with a simplified method based on Equation (6), which models each voxel using only two attributes: signal emission and attenuation. A comparison between the second and third rows in Table 4 shows that our neural approach outperforms the simplified method, highlighting the advantages of voxel-level vector representations and the model’s ability to learn adaptive fusion weights. The difference between the ablated version of GRaF (“without neural ray tracing” in Table 4) and NeRF2 lies in scene representation. While the ablated GRaF retains the parameterized function $\mathcal{T}_{\Psi}$ for processing neighboring spectra, it simplifies voxel modeling and uses the ray tracing module from NeRF2. As shown in Table 2, NeRF2 achieves a PSNR of 17.36 dB, while the ablated version of GRaF yields 16.79 dB (Table 4). This performance gap underscores the importance of GRaF ’s full pipeline.
(a) For AANN training
(b) For AANN testing
Figure 6: Spectrum-based AoA estimation via AANN [2].
5.6 Impact of RF Frequency Bands
To examine the impact of frequency bands, we collect three datasets in a conference room (Figure 4(a)) at 928 MHz, 2.412 GHz, and 5.805 GHz. GRaF is trained on each dataset. As shown in Table 5, GRaF adapts well to different bands while maintaining high-quality synthesized spectra. This design is justified since each band supports different wireless technologies (e.g., 5G, WiFi), requiring dedicated hardware for signal reception and processing [35, 44, 11].
Additionally, in Table 6, we evaluate a version of GRaF trained only on the 2.412 GHz dataset (voxel size $\lambda\approx 0.124\,\text{m}$) and directly tested on the 928 MHz and 5.805 GHz datasets without retraining. During testing, the voxel size is held fixed at $0.124\,\text{m}$. This cross-frequency evaluation reveals a noticeable performance drop, primarily due to two factors. First, RF propagation characteristics vary with frequency. At 928 MHz, signals experience stronger penetration and reduced attenuation, while at 5.805 GHz, they are more prone to scattering and absorption. These variations in wave behavior limit the model’s ability to generalize across bands. Second, maintaining a fixed voxel size leads to spatial resolution mismatch. For 928 MHz ($\lambda\approx 0.323\,\text{m}$), the voxel grid is unnecessarily fine, preserving details but increasing computational cost. In contrast, for 5.805 GHz ($\lambda\approx 0.052\,\text{m}$), the voxel size is too coarse to resolve fine-scale propagation effects.
5.7 Case Study: Angle of Arrival (AoA) Estimation
The synthesized spatial spectrum can be used for downstream tasks such as Angle of Arrival (AoA) estimation for wireless localization [2]. We use an angular artificial neural network (AANN), which consists of a ResNet-50 backbone followed by an MLP head, to process the spectrum and estimate the AoA. This estimated AoA corresponds to the line-of-sight (LoS) propagation direction between the transmitter and receiver, enabling transmitter localization. By generating synthetic training datasets, GRaF significantly reduces the data collection effort required to train the AANN. To demonstrate the benefits of the high-quality spectra synthesized by GRaF, we evaluate using the following two strategies for the training and testing stages:
Table 5: Training separately on each frequency band.
928 MHz 2.412 GHz 5.805 GHz
[…]
Table 6: Training only on the 2.412 GHz frequency band.
[…]
928 MHz 0.104 0.362 14.84 0.576
5.805 GHz 0.218 0.467 11.61 0.493
(i) Synthesized Spectra for AANN Training. An AANN is trained on four datasets: (1) 50% ground truth (GT), (2) 50% GT combined with 50% NeRF2-synthesized spectra, (3) 50% GT combined with 50% GRaF-synthesized spectra, and (4) 100% GT. All training datasets use the same transmitter locations, and each trained AANN is evaluated on a shared testing set with consistent transmitter positions. Figure 6(a) presents the cumulative distribution function (CDF) of the estimated AoA error. The GRaF-augmented dataset reduces the estimation error by 61.6% compared to training on 50% GT alone, and by 25.8% relative to the NeRF2-augmented dataset. These results demonstrate that GRaF produces high-quality spectra that effectively improve localization model training.
(ii) Synthesized Spectra for AANN Testing. An AANN is trained on GT spectra and evaluated on three types of test datasets: GT, NeRF2-synthesized spectra, and GRaF-synthesized spectra. All test spectra correspond to the same transmitter locations. Figure 6(b) demonstrates that GRaF’s spectra yield more accurate AoA estimations, highlighting the superior fidelity of GRaF’s synthesized spectra.
6 Conclusion
This paper presents GRaF, a generalizable RF radiance field framework for spatial spectrum synthesis. We first establish the interpolation theory in the RF domain as the theoretical foundation. GRaF integrates a geometry-aware latent RF radiance field to capture spatially correlated propagation characteristics and employs a neural ray tracing algorithm to aggregate these features for accurate spectrum computation. Experimental results show that GRaF generalizes across scenes and enhances downstream task performance.
Acknowledgements
This research was funded in part by the National Science Foundation through grants CNS-2239458 and CCSS-2525613. Any opinions, findings, and conclusions expressed in this material are those of the authors and do not necessarily reflect the views of the funding agencies.